CRY+VIS Interface Meeting
Date: 2021/6/18 13:00-13:40
Participants: Takahashi, Ikeda, Miyo, Ushiba, Uchiyama, Sato, Yokozawa, Tomaru, Aso、Hirata
- IX: H4 geophone was replaced. Top chamber was closed again. TF measurement of MN is on going.
- IY: Schedule: Rough alignment of TM was scheduled on 6/22.
- EX: Adjustment of Yaw FR and IP tuning was finished. Broken cable was recovered. Top chamber was closed.
- EY: Assembly of payload has been almost finished. Jig for mirror cleaning will be delivered on 6/21. Measurement of viewport is on going.
- Corrugate tube fixing jig: Delivery and implement were scheduled on 7/12-13.
- TM height setting: Next target is ITMY.
- Commissioning in air: X-arm.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check: Group work is set every Friday.
- PF actuator: After SC.
- Jig for FC painting:
- Separation of rack in 2F: On the maintenance days in July.
- Disassembly of PRM/PR2 assembly frame: This was scheduled on 6/28-29.
Participants: Takahashi, Hirata, Sato, Fabian, Tanaka, Node, Aso, Terrence
This is generated by python script in /users/Miyo/dropbox/Script/scrapping/main.py.
- Tuned the F0 in IY.
- Adjusted the yaw FR in EX. Tuned the IP. The H4 geophone was put on the IP. Repaired some in-vacuum cables.
- Closed the top chambers in IX and EX.
Limit switches in SR2 and SR3
There were three switches that didn't seem to be working. Hirata-san fixed them and we checked them with the multimeter. They are working now
- We measured IP TF on Monday:
- The resonant frequencies in L and T looked low enough (63 and 55 mHz) so we didn't adjust the ballast masses.
- The coherence wasn't bad but not excelent either; the frequency resolution was 10 mHz.
- Terrence measured TFs again with a better frequency resolution and found that L and T are highly coupled.
- Fabian locked the SRM IP so I can measure the LVDT sensor noises.
- I measured the geophone readouts as well when the IP is unlocked. The suspension is quiet enough so these can be used as sensor noise estimates. [notebook sneak peek]
- I then measured transfer functions and cross-transfer functions.
cross-coupling is confusing. Explain with figure: klog 17137
- I reverted the sensing matrix and the actuation matrix using Mark's old matrix, which is derived from geometry.
- Cross-coupling is still confusing between L and T. But at least I can decouple L2Y, T2Y, Y2L and Y2T.
- I suspect the lowest frequency modes are L-T coupled. I am pretty sure L and T sensors are not 100% coupled as suggested in the figures.
- I compared the L and T readout with seismic noise in the x-arm and y-arm direction and L has minimal coherence with x-arm while T has minimal coherence with y-arm.
- Peaks are too close :*(.
- Some opinion about the RT-model
- LVDTALIGN and ACCALIGN matrices are great for sensor diagonalization.
- Why don't add something like COILALIGN (for actuation)?
We revisited the BS IP to check stability and values of resonance frequencies.
- When we adjusted the ballast mass before the temperature 22.7 degrees, this week was 23.5 degrees; O3GK temperature is 23 degrees.
I moved the IP close to the setpoints using the horizontal fishing rods: it looked unstable, it began moving by small amount but it suddendly moved a lot (e.g. from -1.5 mm to 1.5 mm). With some care I was able to move it at the setpoint and measured TFs, the resonant frequencies were 43 and 86 mHz for L and T respectively.
We removed 4.5 kg and and the frequencies changed to 93 and 70 mHz for L and T respectively.
- With the same amount of mass we tested different distributions to check whether we can reduce the assymetry between the resonant frequencies in L and T:
- See klog.
The assymetry remains around 25 mHz: one is typically around 65 mHz and the other one around 90 mHz.
- Becuase of temperature changes one might be too low and the other too high.
- L and T seem strongly coupled.
- I wonder how to proceed: during installation we were not able to distribute the masses evenly around the circumference becuase the magic wand (yet to be conforimed with a klog entry).
- Type-A team: ACC preparation
- Type-B team: Remaining work
IMC suspensions are ready in air  will finish 2021/04/21.
PRs are ready in air  will finish 2021/07/09.
ETMX is ready in air  will finish 2021/08/25.
ITMX is ready in air  will finish 2021/08/31.
ETMY is ready in air  will finish 2021/09/16.
ITMY is ready in air  will finish 2021/11/08.
BS is ready in air  will finish 2021/11/24.
SRs are ready in air  will finish 2021/12/20.
Output suspensions are ready in air  will finish 2021/12/21.
TMSY is ready in air  will finish 2021/??/??.
TMSX is ready in air  will finish 2021/??/??.
The week from 6/21
- Takahashi: 6/21-22
- Hirata: 6/21-22
- Yano: 6/21-25
The week from 6/28
- Takahashi: 6/28-7/2
- Sato: 6/28-7/2
- Hirata: 6/28-7/2
- Yano: 6/28-7/2